Solve the equation. 3 7 9) 1 2 3 = 1 + Alw


please help me answer question 10
Answer #1

5-2 + 3 - - 5-a- - - t 7 = -3 -3 4 taking L.Com. in LHS =) 5-7-a = 4 (-2-a) = -3(a) 7 - 8 - 49 = -3Q * - 8 = 40-39 + - 8 = ai have tried my best to solve this problem but still if you have any doubt regarding this question then please comment